ORTHOGONAL POLYNOMIALS AND

MARKOV CHAINS

Erik A. van Doorn^{⋆} and Pauline Schrijner^{⋆⋆}

⋆ Faculty of Applied Mathematics University of Twente

P.O. Box 217, 7500 AE Enschede, The Netherlands

⋆⋆ Department of Mathematics and Statistics

University of Newcastle Newcastle upon Tyne NE1 7RU, England

Keywords and phrases: birth-death process, quasi-stationary distribution, randomwalk, random-walk polynomials, random-walk measure 1991 Mathematics Subject Classification: 42 C 05, 60 J 10